OpenStudy (anonymous):

HELP PLEASE! :( Which of the points is a solution to 2x – 4y = 6? A. (-2, 1) B. (1, -1) C. (1, 1) D. (2, 1)

Directrix (directrix):

Take the equation and test the options. To help you get started, I'll test option A to see if it is a solution.

Directrix (directrix):

Test (-2,1) 2x – 4y = 6 Replace x in the equation by -2 and replace y by 1. Simplify the left side. If it equals 6, the the point is a solution. 2*(-2) – 4*(1) = -4 -4 = -8 which is not 6. The point (-2,1) is not a solution of the equation. Test option B. @lizette_macias11